Calculate the capacitive reactance of 5 microfarad capacitor for a frequency of 50Hz and 10^6 Hz

Grade:12

1 Answers

Aafiya Shaikh
6 years ago
Capacitive reactance is Xc=1/wc,,so,,for 50Hz & 5UF capacitor,Xc=1/5×10^-6×50. = 4×10^3ohm..&. For 10^6Hz & 5UF..,Xc = 1/5×10^-6×10^6 = 0.2ohm